# Large Sets of Asymptotically Optimal and Near-Optimal Quasi-Complementary

### v1

@@ II-C Gaussian sums and character sums over finite fields
    $f$ has exactly one root in ${\mathbb{F}}_{q}$ if and only if $b=0$;
@@
    $f$ has exactly two roots in ${\mathbb{F}}_{q}$ if and only if $b\neq 0$ and ${\mathrm{Tr}}_{q/2}(\frac{ac}{b^{2}})=0$;
@@
    $f$ has no root in ${\mathbb{F}}_{q}$ if and only if $b\neq 0$ and ${\mathrm{Tr}}_{q/2}(\frac{ac}{b^{2}})=1$.
@@
For $a\neq 0$, we substitute $x$ by $x-a$ in $f(x)$. Let $\alpha_{1}:=b-a^{p^{k}}$ and $\beta_{1}:=c-ab$. Then $f(x)$ and be reduced to $g(x)=x^{p^{k}+1}+\alpha_{1}x+\beta_{1}$. If we substitute $x$ by $ux$ in $g(x)$ for $u^{p^{k}}=\alpha$ and $\alpha=b-a^{p^{k}}\neq 0$, then $g(x)$ can be transformed into the form $P_{\gamma}(x)=x^{p^{k}+1}+x+\gamma$, where $\gamma=\frac{\beta}{u^{p^{k}+1}}$. By the results in [1], we can easily derive the number of zeros of $f(x)$.
Lemma II.5
@@
    $N_{f}=\left\{\begin{array}[]{ll}1&\text{if}~{}\frac{n}{h}~{}\text{is}~{}\text{odd}~{}\text{and}~{}p=2,\\ 0,1~{}\text{or}~{}2&\text{if}~{}\frac{n}{h}~{}\text{is}~{}\text{odd}~{}\text{and}~{}p~{}\text{is}~{}\text{odd},\\ 0,1~{}\text{or}~{}p^{h}+1&\text{if}~{}\frac{n}{h}~{}\text{is}~{}\text{even}.\end{array}\right.$
@@ III New constructions of Periodic QCSSs
In this section, we will present several new constructions of asymptotically optimal or near-optimal periodic QCSSs with large set sizes from special polynomials or Gaussian sums.
@@ III-A The first construction of periodic QCSSs from quadratic polynomials
$\mathbf{c}_{l}^{a,b}=\left(\mathbf{c}_{l}^{a,b}(t)\right)_{t=0}^{q-2},\mbox{ where}\ \mathbf{c}_{l}^{a,b}(t)=\chi_{1}(\alpha^{t}f(d_{l}))\mbox{for}~{}0\leq l\leq q-1.$
@@
$\mathcal{C}=\{\mathbf{C}^{a,b}:a\in{\mathbb{F}}_{q},b\in{\mathbb{F}}_{q}\}.$ (9)
@@
where $a_{1},b_{1},a_{2},b_{2}\in{\mathbb{F}}_{q}$. We now consider the following cases to determine the value of $R_{\mathbf{C}^{a_{1},b_{1}},\mathbf{C}^{a_{2},b_{2}}}(\tau)$.
@@
where $G(\eta,\chi_{1})=(-1)^{n-1}(\sqrt{-1})^{(\frac{p-1}{2})^{2}n}\sqrt{q}$ by Lemma II.1. Thus, $|R_{\mathbf{C}^{a_{1},b_{1}},\mathbf{C}^{a_{2},b_{2}}}(\tau)|\in\{0,q\}$. ∎
Summarizing the above four cases, we derive that the maximum periodic correlation magnitude of $\mathcal{C}$ is $q$. Next, we will show that the parameters of the obtained periodic QCSS asymptotically achieving the correlation lower bound in (1). Since $\mathcal{C}$ is a periodic $(q^{2},q,q-1,q)$-QCSS, according to the bound in (1), we have
$\vartheta_{opt}=q(q-1)\sqrt{\frac{q-1}{q^{2}(q-1)-1)}}=q\sqrt{\frac{(q-1)^{3}}{q^{2}(q-1)-1)}}.$
@@
$\lim_{q\rightarrow+\infty}\frac{\vartheta_{max}}{\vartheta_{opt}}=\lim_{q\rightarrow+\infty}\frac{q}{q\sqrt{\frac{(q-1)^{3}}{q^{2}(q-1)-1)}}}=1.$
@@
where $a_{1},b_{1},a_{2},b_{2}\in{\mathbb{F}}_{q}$. Then we consider the following cases to determine the value of $R_{\mathbf{C}^{a_{1},b_{1}},\mathbf{C}^{a_{2},b_{2}}}(\tau)$.
@@
$R_{\mathbf{C}^{a_{1},b_{1}},\mathbf{C}^{a_{2},b_{2}}}(\tau)=\left\{\begin{array}[]{ll}-q&\text{if}~{}a_{1}=a_{2},b_{1}\neq b_{2},\\ 0&\text{if}~{}a_{1}\neq a_{2},b_{1}=b_{2},\\ 0&\text{if}~{}a_{1}\neq a_{2},b_{1}\neq b_{2}.\end{array}\right.$
@@
By LemmaII.4, we know that $f(x)=(1-\alpha^{\tau})x^{2}+b_{1}-b_{2}\alpha^{\tau}$ has exactly one root in ${\mathbb{F}}_{q}$. By the orthogonal relation of the additive characters, we have
@@
$=$ $\sum_{x\in{\mathbb{F}}_{q}}\sum_{y\in{\mathbb{F}}_{q}^{*}}\chi_{1}((1-\alpha^{\tau})yx^{2}+(a_{1}-a_{2}\alpha^{\tau})yx+(b_{1}-b_{2}\alpha^{\tau})y)$
$=$ $\sum_{x\in{\mathbb{F}}_{q}}\sum_{y\in{\mathbb{F}}_{q}^{*}}\chi_{1}(((1-\alpha^{\tau})x^{2}+(a_{1}-a_{2}\alpha^{\tau})x+(b_{1}-b_{2}\alpha^{\tau}))y).$
By Lemma II.4, we know that $f(x)=(1-\alpha^{\tau})x^{2}+(a_{1}-a_{2}\alpha^{\tau})x+(b_{1}-b_{2}\alpha^{\tau})$ has exactly two roots if ${\mathrm{Tr}}_{q/2}\left(\frac{(1-\alpha^{\tau})(b_{1}-b_{2}\alpha^{\tau})}{(a_{1}-a_{2}\alpha^{\tau})^{2}}\right)=0$, and no root if ${\mathrm{Tr}}_{q/2}\left(\frac{(1-\alpha^{\tau})(b_{1}-b_{2}\alpha^{\tau})}{(a_{1}-a_{2}\alpha^{\tau})^{2}}\right)=1$. By the orthogonal relation of the additive characters, we have
$R_{\mathbf{C}^{a_{1},b_{1}},\mathbf{C}^{a_{2},b_{2}}}(\tau)=\left\{\begin{array}[]{ll}-q&\text{if}~{}{\mathrm{Tr}}_{q/2}\left(\frac{(1-\alpha^{\tau})(b_{1}-b_{2}\alpha^{\tau})}{(a_{1}-a_{2}\alpha^{\tau})^{2}}\right)=1,\\ q&\text{if}~{}{\mathrm{Tr}}_{q/2}\left(\frac{(1-\alpha^{\tau})(b_{1}-b_{2}\alpha^{\tau})}{(a_{1}-a_{2}\alpha^{\tau})^{2}}\right)=0.\end{array}\right.$
@@
Example 2
[…]
@@ III-B The second construction of periodic QCSSs from a special polynomial
$\mathbf{f}_{l}^{a,b,c}=\left(\mathbf{f}_{l}^{a,b,c}(t)\right)_{t=0}^{q-2},~{}~{}\mbox{where }\mathbf{f}_{l}^{a,b,c}(t)=\chi_{1}(\alpha^{t}f(d_{l})),\mbox{ for }0\leq l\leq q-1.$ (14)
@@
=$\mathcal{F}=\{\mathbf{F}^{a,b,c}:a\in{\mathbb{F}}_{q},b\in{\mathbb{F}}_{q},c\in{\mathbb{F}}_{q}\}.$ (16)
Theorem III.3
@@
We divide into the following cases to determine the value of $R_{\mathbf{F}^{a_{1},b_{1},c_{1}},\mathbf{F}^{a_{2},b_{2},c_{2}}}(\tau)$.
@@
$=$ $\left\{\begin{array}[]{ll}-q&\text{if}~{}b_{1}=b_{2}~{}\text{and}~{}c_{1}\neq c_{2},\\ 0&\text{if}~{}b_{1}\neq b_{2}~{}\text{and}~{}c_{1}=c_{2},\\ 0&\text{if}~{}b_{1}\neq b_{2}~{}\text{and}~{}c_{1}\neq c_{2}.\end{array}\right.$
@@
$=$ $\left\{\begin{array}[]{ll}-q&\text{if}~{}N_{g}=0,\\ 0&\text{if}~{}N_{g}=1,\\ q&\text{if}~{}N_{g}=2.\end{array}\right.$
@@
$=$ $\left\{\begin{array}[]{ll}-q&\text{if}~{}N_{f}=0,\\ 0&\text{if}~{}N_{f}=1,\\ q&\text{if}~{}N_{f}=2,\\ 2q&\text{if}~{}N_{f}=3.\end{array}\right.$
@@
$\vartheta_{opt}=q(q-1)\sqrt{\frac{q^{2}-1}{q^{3}(q-1)-1}}=q\sqrt{\frac{(q-1)^{3}(q+1)}{q^{3}(q-1)-1}}.$
@@
$\lim_{q\rightarrow+\infty}\frac{\vartheta_{max}}{\vartheta_{opt}}=\lim_{q\rightarrow+\infty}\frac{2q}{q\sqrt{\frac{(q-1)^{3}(q+1)}{q^{3}(q-1)-1)}}}=2.$
@@ III-C The third construction of periodic QCSSs from a cubic polynomial
$\mathbf{h}_{l}^{a,b,c}=\left(\mathbf{h}_{l}^{a,b,c}(t)\right)_{t=0}^{q-2},~{}~{}\text{where}\ \mathbf{h}_{l}^{a,b,c}(t)=\chi_{1}\left(\alpha^{t}f(d_{l})\right)\mbox{ for }0\leq l\leq q-1.$
@@
=$\mathcal{H}=\{\mathbf{H}^{a,b,c}:a\in{\mathbb{F}}_{q},b\in{\mathbb{F}}_{q},c\in{\mathbb{F}}_{q}\}.$ (30)
Theorem III.4
@@
where $a_{1},a_{2},b_{1},b_{2},c_{1},c_{2}\in{\mathbb{F}}_{q}$. We consider the following cases to determine its value.
@@
Now we consider the zeros of polynomial $g(x)=(a_{1}-a_{2})x^{2}+(b_{1}-b_{2})x+c_{1}-c_{2}$ in ${\mathbb{F}}_{q}$ with $a_{1}-a_{2}\neq 0$, where $p$ is an odd prime. Let $N_{g}$ denote the number of roots of $g(x)=0$ in ${\mathbb{F}}_{q}$. It is easy to know that $N_{g}\in\{0,1,2\}$. By the orthogonal relation of the additive characters, we have
@@
Now we consider the zeros of the polynomial $f(x)=(1-\alpha^{\tau})x^{3}+(a_{1}-a_{2}\alpha^{\tau})x^{2}+(b_{1}-b_{2}\alpha^{\tau})x+c_{1}-c_{2}\alpha^{\tau}$ in ${\mathbb{F}}_{q}$ with $1-\alpha^{\tau}\neq 0$, where $p$ is an odd prime. Let $N_{f}$ denote the number of zeros of $f$ in ${\mathbb{F}}_{q}$. It is known that $N_{f}\in\{0,1,2,3\}$. Note that the each of the coefficients $a_{1}-a_{2}\alpha^{\tau}$, $b_{1}-b_{2}\alpha^{\tau}$ and $c_{1}-c_{2}\alpha^{\tau}$ runs over ${\mathbb{F}}_{q}$ when all of $a_{1},a_{1},b_{1},b_{2},c_{1},c_{2}$ runs over ${\mathbb{F}}_{q}$. Note that there exits a polynomial $f(x)=(1-\alpha^{\tau})(x-x_{1})(x-x_{2})(x-x_{3})$ for pairwise distinct elements $x_{1},x_{2},x_{3}\in{\mathbb{F}}_{q}$ which has three zeros in ${\mathbb{F}}_{q}$. By the orthogonal relation of additive characters, we then have
@@
$\vartheta_{opt}=q(q-1)\sqrt{\frac{q^{2}-1}{q^{3}(q-1)-1}}=q\sqrt{\frac{(q-1)^{3}(q+1)}{q^{3}(q-1)-1}}.$
@@
$\lim_{q\rightarrow+\infty}\frac{\vartheta_{max}}{\vartheta_{opt}}=\lim_{q\rightarrow+\infty}\frac{2q}{q\sqrt{\frac{(q-1)^{3}(q+1)}{q^{3}(q-1)-1)}}}=2.$
@@
Example 4
Let $p=3$ and $n=2$. Then the parameters of the QCSS $\mathcal{H}$ constructed in Theorem III.4 is $(729,9,8,18)$ and its alphabet is given by $\{e^{2\pi\sqrt{-1}i/3}:i\in[0,2]\}$. By Magma program, we verify that the periodic correlation function $R_{\mathbf{H}^{a_{1},b_{1},c_{1}},\mathbf{H}^{a_{2},b_{2},c_{1}}}(\tau)$ is equal to 00, $9$ or $18$ for any $a,b,c\in{\mathbb{F}}_{q}$ and $0\leq\tau\leq 7$ except the trivial case that $a_{1}=a_{2},b_{1}=b_{2},c_{1}=c_{2}$ and $\tau=0$. For instance, the matrices $\mathbf{H}^{1,\alpha,\alpha}$ and $\mathbf{H}^{1,\alpha,\alpha^{2}}$ in $\mathcal{H}$ are presented as follows, where each entry stands for a power of $\zeta_{3}=e^{2\pi\sqrt{-1}/3}$. By Python program, we show the autocorrelation magnitude distribution of $\mathbf{H}^{1,\alpha,\alpha}$ in Fig. $10$, the autocorrelation magnitude distribution of $\mathbf{H}^{1,\alpha,\alpha^{2}}$ in Fig. $11$, and the correlation magnitude distribution of $\mathbf{H}^{1,\alpha,\alpha}$ and $\mathbf{H}^{1,\alpha,\alpha^{2}}$ in Fig. $12$. We remark that the upper bound of $\vartheta_{max}$ in Theorem III.4 is tight in this case.

### v2

@@ II-D The number of zeros of some polynomials over finite fields
    $f$ has exactly one zero in ${\mathbb{F}}_{q}$ if and only if $b=0$;
@@
    $f$ has exactly two zeros in ${\mathbb{F}}_{q}$ if and only if $b\neq 0$ and ${\mathrm{Tr}}_{q/2}(\frac{ac}{b^{2}})=0$;
@@
    $f$ has no zero in ${\mathbb{F}}_{q}$ if and only if $b\neq 0$ and ${\mathrm{Tr}}_{q/2}(\frac{ac}{b^{2}})=1$.
@@
For $a\neq 0$, we substitute $x$ by $x-a$ in $f(x)$. Let $\alpha_{1}:=b-a^{p^{k}}$ and $\beta_{1}:=c-ab$. Then $f(x)$ can be reduced to $g(x)=x^{p^{k}+1}+\alpha_{1}x+\beta_{1}$. If we substitute $x$ by $ux$ in $g(x)$ for $u^{p^{k}}=\alpha$ and $\alpha=b-a^{p^{k}}\neq 0$, then $g(x)$ can be transformed into the form $P_{\gamma}(x)=x^{p^{k}+1}+x+\gamma$, where $\gamma=\frac{\beta}{u^{p^{k}+1}}$. By the results in [1], we can easily derive the number of zeros of $f(x)$ in ${\mathbb{F}}_{q}$.
Lemma II.5.
@@
    $N_{f}=\left\{\begin{array}[]{ll}1&\text{if}\leavevmode\nobreak\ \frac{n}{h}\leavevmode\nobreak\ \text{is}\leavevmode\nobreak\ \text{odd}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ p=2,\\ 0,1\leavevmode\nobreak\ \text{or}\leavevmode\nobreak\ 2&\text{if}\leavevmode\nobreak\ \frac{n}{h}\leavevmode\nobreak\ \text{is}\leavevmode\nobreak\ \text{odd}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ p\leavevmode\nobreak\ \text{is}\leavevmode\nobreak\ \text{odd},\\ 0,1\leavevmode\nobreak\ \text{or}\leavevmode\nobreak\ p^{h}+1&\text{if}\leavevmode\nobreak\ \frac{n}{h}\leavevmode\nobreak\ \text{is}\leavevmode\nobreak\ \text{even}.\end{array}\right.$
@@ III New constructions of periodic QCSSs
In this section, we will present several new constructions of asymptotically optimal or near-optimal periodic QCSSs with large set sizes from special polynomials over finite fields or Gaussian sums.
@@ III-A The first construction of periodic QCSSs from quadratic polynomials
$\mathbf{c}_{l}^{a,b}=\left(\mathbf{c}_{l}^{a,b}(t)\right)_{t=0}^{q-2},\mbox{ where}\ \mathbf{c}_{l}^{a,b}(t)=\chi_{1}(\alpha^{t}f(d_{l}))\mbox{for}\leavevmode\nobreak\ 0\leq l\leq q-1.$
@@
$\mathcal{C}=\left\{\mathbf{C}^{a,b}:a\in{\mathbb{F}}_{q},b\in{\mathbb{F}}_{q}\right\}.$ (9)
@@
where $a_{1},b_{1},a_{2},b_{2}\in{\mathbb{F}}_{q}$. We now consider the following cases to determine the value distribution of $R_{\mathbf{C}^{a_{1},b_{1}},\mathbf{C}^{a_{2},b_{2}}}(\tau)$.
@@
where $G(\eta,\chi_{1})=(-1)^{n-1}(\sqrt{-1})^{(\frac{p-1}{2})^{2}n}\sqrt{q}$ by Lemma II.1. Thus, $|R_{\mathbf{C}^{a_{1},b_{1}},\mathbf{C}^{a_{2},b_{2}}}(\tau)|\in\{0,q\}$.
Summarizing the above four cases, we derive that the maximum periodic correlation magnitude of $\mathcal{C}$ is $q$. Next, we will show that the parameters of the obtained periodic QCSS asymptotically achieve the correlation lower bound in (1). Since $\mathcal{C}$ is a periodic $(q^{2},q,q-1,q)$-QCSS, according to the bound in (1), we have
$\vartheta_{{\text{opt}}}=q(q-1)\sqrt{\frac{q-1}{q^{2}(q-1)-1)}}=q\sqrt{\frac{(q-1)^{3}}{q^{2}(q-1)-1)}}.$
@@
$\lim_{q\rightarrow+\infty}\frac{\vartheta_{\max}}{\vartheta_{{\text{opt}}}}=\lim_{q\rightarrow+\infty}\frac{q}{q\sqrt{\frac{(q-1)^{3}}{q^{2}(q-1)-1)}}}=1.$
@@
where $a_{1},b_{1},a_{2},b_{2}\in{\mathbb{F}}_{q}$. Then we consider the following cases to determine the value distribution of $R_{\mathbf{C}^{a_{1},b_{1}},\mathbf{C}^{a_{2},b_{2}}}(\tau)$.
@@
$R_{\mathbf{C}^{a_{1},b_{1}},\mathbf{C}^{a_{2},b_{2}}}(\tau)=\left\{\begin{array}[]{ll}-q&\text{if}\leavevmode\nobreak\ a_{1}=a_{2},b_{1}\neq b_{2},\\ 0&\text{if}\leavevmode\nobreak\ a_{1}\neq a_{2},b_{1}=b_{2},\\ 0&\text{if}\leavevmode\nobreak\ a_{1}\neq a_{2},b_{1}\neq b_{2}.\end{array}\right.$
@@
By Lemma II.4, we know that $f(x)=(1-\alpha^{\tau})x^{2}+b_{1}-b_{2}\alpha^{\tau}$ has exactly one zero in ${\mathbb{F}}_{q}$. By the orthogonal relation of the additive characters, we have
@@
$=$ $\sum_{x\in{\mathbb{F}}_{q}}\sum_{y\in{\mathbb{F}}_{q}^{*}}\chi_{1}\left((1-\alpha^{\tau})yx^{2}+(a_{1}-a_{2}\alpha^{\tau})yx+(b_{1}-b_{2}\alpha^{\tau})y\right)$
$=$ $\sum_{x\in{\mathbb{F}}_{q}}\sum_{y\in{\mathbb{F}}_{q}^{*}}\chi_{1}\left(\left((1-\alpha^{\tau})x^{2}+(a_{1}-a_{2}\alpha^{\tau})x+(b_{1}-b_{2}\alpha^{\tau})\right)y\right).$
By Lemma II.4, we know that $f(x)=(1-\alpha^{\tau})x^{2}+(a_{1}-a_{2}\alpha^{\tau})x+(b_{1}-b_{2}\alpha^{\tau})$ has exactly two zeros if ${\mathrm{Tr}}_{q/2}\left(\frac{(1-\alpha^{\tau})(b_{1}-b_{2}\alpha^{\tau})}{(a_{1}-a_{2}\alpha^{\tau})^{2}}\right)=0$, and no zero if ${\mathrm{Tr}}_{q/2}\left(\frac{(1-\alpha^{\tau})(b_{1}-b_{2}\alpha^{\tau})}{(a_{1}-a_{2}\alpha^{\tau})^{2}}\right)=1$. By the orthogonal relation of the additive characters, we have
$R_{\mathbf{C}^{a_{1},b_{1}},\mathbf{C}^{a_{2},b_{2}}}(\tau)=\left\{\begin{array}[]{ll}-q&\text{if}\leavevmode\nobreak\ {\mathrm{Tr}}_{q/2}\left(\frac{(1-\alpha^{\tau})(b_{1}-b_{2}\alpha^{\tau})}{(a_{1}-a_{2}\alpha^{\tau})^{2}}\right)=1,\\ q&\text{if}\leavevmode\nobreak\ {\mathrm{Tr}}_{q/2}\left(\frac{(1-\alpha^{\tau})(b_{1}-b_{2}\alpha^{\tau})}{(a_{1}-a_{2}\alpha^{\tau})^{2}}\right)=0.\end{array}\right.$
@@
Example 2.
Let $p=2$ and $n=4$. Then the parameters of the QCSS $\mathcal{C}$ constructed in Theorem III.2 are $(256,16,15,16)$ and its alphabet is given by $\{(-1)^{i}:i\in[0,1]\}$. By Magma program, we verify that the periodic correlation function $R_{\mathbf{C}^{a_{1},b_{1}},\mathbf{C}^{a_{2},b_{2}}}(\tau)$ is equal to 00 or $16$ for any $a_{1},b_{1},a_{2},b_{2}\in{\mathbb{F}}_{q}$ and $0\leq\tau\leq 14$ except the trivial case that $a_{1}=a_{2},b_{1}=b_{2}$ and $\tau=0$. For instance, the matrices $\mathbf{C}^{1,\alpha}$ and $\mathbf{C}^{1,\alpha^{2}}$ in $\mathcal{C}$ are presented as follows, where each entry stands for a power of $\zeta_{2}=-1$. By Python program, we show the autocorrelation magnitude distribution of $\mathbf{C}^{1,\alpha}$ in Fig. 4, the autocorrelation magnitude distribution of $\mathbf{C}^{1,\alpha^{2}}$ in Fig. 5, and the correlation magnitude distribution of $\mathbf{C}^{1,\alpha}$ and $\mathbf{C}^{1,\alpha^{2}}$ in Fig. 6, respectively.
@@ III-B The second construction of periodic QCSSs from a special polynomial
$\mathbf{f}_{l}^{a,b,c}=\left(\mathbf{f}_{l}^{a,b,c}(t)\right)_{t=0}^{q-2},\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{where }\mathbf{f}_{l}^{a,b,c}(t)=\chi_{1}(\alpha^{t}f(d_{l}))\mbox{ for }0\leq l\leq q-1.$ (14)
@@
=$\mathcal{F}=\left\{\mathbf{F}^{a,b,c}:a\in{\mathbb{F}}_{q},b\in{\mathbb{F}}_{q},c\in{\mathbb{F}}_{q}\right\}.$ (16)
Theorem III.3.
@@
We divide into the following cases to determine the value distribution of $R_{\mathbf{F}^{a_{1},b_{1},c_{1}},\mathbf{F}^{a_{2},b_{2},c_{2}}}(\tau)$.
@@
$=$ $\left\{\begin{array}[]{ll}-q&\text{if}\leavevmode\nobreak\ b_{1}=b_{2}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ c_{1}\neq c_{2},\\ 0&\text{if}\leavevmode\nobreak\ b_{1}\neq b_{2}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ c_{1}=c_{2},\\ 0&\text{if}\leavevmode\nobreak\ b_{1}\neq b_{2}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ c_{1}\neq c_{2}.\end{array}\right.$
@@
$=$ $\left\{\begin{array}[]{ll}-q&\text{if}\leavevmode\nobreak\ N_{g}=0,\\ 0&\text{if}\leavevmode\nobreak\ N_{g}=1,\\ q&\text{if}\leavevmode\nobreak\ N_{g}=2.\end{array}\right.$
@@
$=$ $\left\{\begin{array}[]{ll}-q&\text{if}\leavevmode\nobreak\ N_{f}=0,\\ 0&\text{if}\leavevmode\nobreak\ N_{f}=1,\\ q&\text{if}\leavevmode\nobreak\ N_{f}=2,\\ 2q&\text{if}\leavevmode\nobreak\ N_{f}=3.\end{array}\right.$
@@
$\vartheta_{{\text{opt}}}=q(q-1)\sqrt{\frac{q^{2}-1}{q^{3}(q-1)-1}}=q\sqrt{\frac{(q-1)^{3}(q+1)}{q^{3}(q-1)-1}}.$
@@
$\lim_{q\rightarrow+\infty}\frac{\vartheta_{\max}}{\vartheta_{{\text{opt}}}}=\lim_{q\rightarrow+\infty}\frac{2q}{q\sqrt{\frac{(q-1)^{3}(q+1)}{q^{3}(q-1)-1)}}}=2.$
@@ III-C The third construction of periodic QCSSs from a cubic polynomial
$\mathbf{h}_{l}^{a,b,c}=\left(\mathbf{h}_{l}^{a,b,c}(t)\right)_{t=0}^{q-2},\leavevmode\nobreak\ \leavevmode\nobreak\ \text{where}\ \mathbf{h}_{l}^{a,b,c}(t)=\chi_{1}\left(\alpha^{t}f(d_{l})\right)\mbox{ for }0\leq l\leq q-1.$
@@
=$\mathcal{H}=\left\{\mathbf{H}^{a,b,c}:a\in{\mathbb{F}}_{q},b\in{\mathbb{F}}_{q},c\in{\mathbb{F}}_{q}\right\}.$ (30)
Theorem III.4.
@@
where $a_{1},a_{2},b_{1},b_{2},c_{1},c_{2}\in{\mathbb{F}}_{q}$. We consider the following cases to determine its value distribution.
@@
Now we consider the zeros of the polynomial $g(x)=(a_{1}-a_{2})x^{2}+(b_{1}-b_{2})x+c_{1}-c_{2}$ in ${\mathbb{F}}_{q}$ with $a_{1}-a_{2}\neq 0$, where $p$ is an odd prime. Let $N_{g}$ denote the number of roots of $g(x)=0$ in ${\mathbb{F}}_{q}$. It is easy to know that $N_{g}\in\{0,1,2\}$. By the orthogonal relation of the additive characters, we have
@@
Now we consider the zeros of the polynomial $f(x)=(1-\alpha^{\tau})x^{3}+(a_{1}-a_{2}\alpha^{\tau})x^{2}+(b_{1}-b_{2}\alpha^{\tau})x+c_{1}-c_{2}\alpha^{\tau}$ in ${\mathbb{F}}_{q}$ with $1-\alpha^{\tau}\neq 0$, where $p$ is an odd prime. Let $N_{f}$ denote the number of zeros of $f$ in ${\mathbb{F}}_{q}$. It is known that $N_{f}\in\{0,1,2,3\}$. Note that the each of the coefficients $a_{1}-a_{2}\alpha^{\tau}$, $b_{1}-b_{2}\alpha^{\tau}$ and $c_{1}-c_{2}\alpha^{\tau}$ runs over ${\mathbb{F}}_{q}$ when all of $a_{1},a_{1},b_{1},b_{2},c_{1},c_{2}$ run over ${\mathbb{F}}_{q}$. Note that there exists a polynomial $f(x)=(1-\alpha^{\tau})(x-x_{1})(x-x_{2})(x-x_{3})$ for pairwise distinct elements $x_{1},x_{2},x_{3}\in{\mathbb{F}}_{q}$ which has three zeros in ${\mathbb{F}}_{q}$. By the orthogonal relation of additive characters, we then have
@@
$\vartheta_{{\text{opt}}}=q(q-1)\sqrt{\frac{q^{2}-1}{q^{3}(q-1)-1}}=q\sqrt{\frac{(q-1)^{3}(q+1)}{q^{3}(q-1)-1}}.$
@@
$\lim_{q\rightarrow+\infty}\frac{\vartheta_{\max}}{\vartheta_{{\text{opt}}}}=\lim_{q\rightarrow+\infty}\frac{2q}{q\sqrt{\frac{(q-1)^{3}(q+1)}{q^{3}(q-1)-1)}}}=2.$
@@
Example 4.
Let $p=3$ and $n=2$. Then the parameters of the QCSS $\mathcal{H}$ constructed in Theorem III.4 are $(729,9,8,18)$ and its alphabet is given by $\{e^{2\pi\sqrt{-1}i/3}:i\in[0,2]\}$. By Magma program, we verify that the periodic correlation function $R_{\mathbf{H}^{a_{1},b_{1},c_{1}},\mathbf{H}^{a_{2},b_{2},c_{1}}}(\tau)$ is equal to 00, $9$ or $18$ for any $a,b,c\in{\mathbb{F}}_{q}$ and $0\leq\tau\leq 7$ except the trivial case that $a_{1}=a_{2},b_{1}=b_{2},c_{1}=c_{2}$ and $\tau=0$. For instance, the matrices $\mathbf{H}^{1,\alpha,\alpha}$ and $\mathbf{H}^{1,\alpha,\alpha^{2}}$ in $\mathcal{H}$ are presented as follows, where each entry stands for a power of $\zeta_{3}=e^{2\pi\sqrt{-1}/3}$. By Python program, we show the autocorrelation magnitude distribution of $\mathbf{H}^{1,\alpha,\alpha}$ in Fig. $10$, the autocorrelation magnitude distribution of $\mathbf{H}^{1,\alpha,\alpha^{2}}$ in Fig. $11$, and the correlation magnitude distribution of $\mathbf{H}^{1,\alpha,\alpha}$ and $\mathbf{H}^{1,\alpha,\alpha^{2}}$ in Fig. $12$. We remark that the upper bound of $\vartheta_{\max}$ in Theorem III.4 is tight in this case.